CAT Chat Minutes

April 24, 1998

Issues that require action and/or follow-up

1. The orbit problem we experienced during the last run, could it have been connected to a power outage, as it occurred right after the lightning storm?

The Machine Physics people have never indicated that the problem was connected to electrical. The problem is in the stability of the beam, primarily the current in the bunches that trigger the particle BPM. They were unstable due to larger current in each of these six bunches.

Information/Follow-up Items

APS Operational Status Update

The shielding verification schedule for the next cycle (1998-3) and the beamline operations schedule are now available on the web. We have not yet released the days set aside for SOMs that will not be used. We will advise you.

In the past two weeks, we have been in a scheduled maintenance period. Maintenance work was performed on all the machines: linac, PAR, and booster. In the storage ring, some work on rf tuners is in progress. Also some of the storage ring sectors were baked.

Beamline 5 ID is being put back together; we will begin baking next week.

Information

Dr. Eastman has decided to pursue his research career at the U of Chicago as a professor. A search for a replacement will be made by The University of Chicago, which operates ANL.

We will be cleaning the masks outside the FOE in the next two weeks. If you require additional information on the process, please contact M. Ramanathan (mohan@aps.anl.gov).

May 2 is the date of the power cutback. We should experience no problems. Please remember to turn off your computers and lights. The cutback will last ~ 8 hours.